OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (anonymous):

Find slope first: y - y / x - x 6 - (-4) = 10 4 - (-1) = -5 slope: 10/5 = 2/1 = 2 so the slope of a parrallel line would be the same: 2

OpenStudy (anonymous):

woops it was supposed to be 5 not -5 though the answer doesn't change; the slope is still 2/1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np:)